# PuzzleBench: Can LLMs Solve Challenging First-Order Combinatorial Reasoning Problems?

### v2

@@ 5.2 Our Approach
  • Prompting LLMs: The LLM is prompted with $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$ to generate an input-agnostic Python program $\psi$. The LLM is instructed to write $\psi$ to read an input from a file, convert it to a symbolic representation according to the inferred specification of the problem, pass the symbolic representation to the solver and then use the solution from the solver to generate the output in the desired format and write it to another file. The LLM is also prompted with information about the solver and its the underlying language in which the symbolic representation is to be generated. Optionally we can also provide the LLM with a subset of $\mathcal{D}_{\mathcal{P}}$.
@@
  • Symbolic Solver: $\psi$ can convert any input instance $x$ to $(E_{x},V_{x})$ which it passes to the symbolic solver. The solver is agnostic to how the representation $(E_{x},V_{x})$ was created and tries to find an assignment $A_{x}$ to $V_{x}$ which satisfies $E_{x}$ which is passed back to $\psi$.
@@
  • Generating the Final Output: $\psi$ then uses $\mathcal{A}_{x}$ to generate the predicted output $\hat{y}$. This step is need because the symbolic representation was created by $\psi$ and it must recover the desired output representation from $\mathcal{A}_{x}$ which might not be straightforward for all problem representations.
5.3 Verification from Solved Examples
@@
We make use of $\mathcal{D}_{\mathcal{P}}$ to verify and correct (if needed) $\psi$. For each $(x,y)\in\mathcal{D}_{\mathcal{P}}$ (solved input-output pair), we run $\psi$ on $x$ to generate the prediction $\hat{y}$, during which the following can happen: 1) Errors during execution of $\psi$; 2) The solver is unable to find $\mathcal{A}_{x}$ under a certain time limit; 3) $\hat{y}\neq y$ i.e. the predicted output is incorrect; 4) $\hat{y}=y$ i.e. the predicted output is correct. If for any training input one of the first three cases occur we provide feedback to the LLM through prompts to improve and generate a new program. This process is repeated till all training examples are solved correctly or till a maximum number of feedback rounds is reached. The feedback is simple in nature and includes the nature of the error, the actual error from the interpreter/symbolic solver and the input instance on which the error was generated. For example, in the case where the output doesn’t match the gold output we prompt the LLM with the solved example it got wrong and the expected solution. In the case where the solver is not able to solve for the constraints in the time limit, we prompt the LLM to optimize the programs. Appendix B contains details of feedback prompts.
@@
It is possible that a single run of Puzzle-LM does not generate the correct solution for all training examples. Puzzle-LM is run multiple times for a given problem. Given the probabilistic nature of LLMs a new program is generated each time and a new feedback process continues. For the final program, we pick the best program generated during these runs, as judged by the accuracy on the training set. Figure  3 describes our entire approach diagrammatically.
Figure 3: Puzzle-LM: Solid lines indicate the main flow and dotted lines indicate feedback pathways.
Random Guessing Few-Shot Prompting PAL Logic-LM Puzzle-LM
Model - + - + - +
Mixtral 8x7B 23.6% 16.1% 37.92% 0.0% 1.61% 10.77% 36.72%
GPT-3.5-Turbo 17.91% 25.4% 35.82% 53.4% 6.87% 7.41% 30.74% 56.86%
GPT-4-Turbo 30.5% 48.51% 66.45% 37.34% 39.91% 60.48% 93.55%
Table 1: Main Results for PuzzleBench and Puzzle-LM. - and + indicate results before and after refinement.
@@ 6 Experimental Setup
Our experiments answer these research questions:
  1. 1.
    How well do LLMs perform on PuzzleBench?
  2. 2.
    How does Puzzle-LM compare with other tool-augmented LLMs on fcore problems?
  3. 3.
    How useful is using feedback from solved examples and multiple runs for fcore problems?
@@
Baselines: We compare our method with 3 baselines: 1) Standard LLM prompting, which leverages in-context learning to directly answer the questions; 2) Program-aided Language Models, which use imperative programs for reasoning and offload the solution step to a program interpreter; 3) Logic-LM, which offloads the reasoning to a symbolic solver. We use Z3 De Moura and Bjørner (2008) an efficient SMT solver for experiments with Logic-LM and Puzzle-LM. We use the Python interpreter for experiments with PAL and Puzzle-LM. We also evaluate refinement for PAL and Puzzle-LM by using 4 rounds of feedback on solved examples and 5 runs for each problem. We evaluate refinement for Logic-LM by providing 4 rounds of feedback to correct syntactic errors in constraints (if any) for each problem instance.
@@
Models: For each method, we experiment with 3 LLMs: GPT-4-Turbo (gpt-4-0125-preview) OpenAI (2023) which is a SOTA LLM by OpenAI, GPT-3.5-Turbo (gpt-3.5-turbo-0125), a relatively smaller LLM by OpenAI and Mixtral 8x7B (mistral-small) Jiang et al. (2024), an open-source mixture-of-experts model developed by Mistral AI. For experiments with few-shot prompting and Logic-LM, we set the temperature to 00 for reproducible results. For experiments with PAL and Puzzle-LM we use a temperature of $0.7$ to sample several runs, each with a potentially different feedback cycle.
@@
Prompting LLMs: For each method the prompt contains the natural language description of the rules of the problem and the input-output format. It also contains two solved examples – pairs of input and one its corresponding outputs – in the prompt. No additional intermediate supervision (e.g., SMT or Python program) is given in the prompt. For few-shot prompting we directly prompt the LLM to solve test set instances (each tested separately). For PAL we prompt the LLM to write an input-agnostic Python program which reads the input from a file, reasons to solve the input and then writes the solution to another file, the program generated is run on each testing set instance. For Logic-LM for each test set instance we prompt the LLM to convert it into its symbolic representation which is then fed to a symbolic solver, the prompt additionally contains the description of the language of the solver. We then prompt the LLM with the solution from the solver and ask it to generate the output in the desired format. Details for Puzzle-LM can be found in Section 5.1. Prompt templates can be found in Appendix B and other experimental details can be found in Appendix C.
@@
Metrics: For each problem we use the verification script to check if the candidate solution generated for each test set instance is correct. We use this script to compute the accuracy, i.e. fraction of test set instances solved correctly for each problem using binary marking i.e. 0 for incorrect solution and 1 for a correct solution. It is possible that no solution is generated or the solution generated is not in the output-format specified, we consider these to be incorrect and assign a score of 0. We use the macro-average of test set accuracies of all problems in PuzzleBench as our metric.
@@ 7 Results
Table 1 describes the main results for Puzzle-LM and other baselines on PuzzleBench. GPT-4-Turbo is significantly better than other models across all methods. Mixtral 8x7B struggles on our benchmark indicating that smaller LLMs (even with mixture of experts) are not as effective at complex reasoning tasks. Mixtral’s performance with few-shot prompting is slightly better than random guessing and with other methods (without refinement) is even worse than random guessing. PAL and Puzzle-LM tend to perform better than other baselines benefiting from the vast pre-training of LLMs on code Brown et al. (2020); Chen et al. (2021). Logic-LM performs rather poorly with smaller LLMs indicating that they struggle to invoke symbolic solvers directly. Hereafter, we focus on just GPT-4-Turbo’s performance, since it is far superior to other models.
(a) Effect of Feedback
(b) Effect of Mutliple Runs
Figure 4: Dashed lines show results for individual problems in PuzzleBench, with coloured lines highlighting specific cases and grey lines others. The red bold represents the average effect across all problems with GPT-4-Turbo.
7.1 Puzzle-LM vs Baselines
Puzzle-LM outperforms few-shot prompting and Logic-LM across all problems in PuzzleBench. On average with GPT-4-Turbo the improvements are by an impressive $63.05\%$ against few-shot prompting and by $53.64\%$ against Logic-LM on PuzzleBench (with refinement). Few-shot prompting struggles, solving less than a third of the problems even with GPT-4-Turbo indicating the need for augmenting LLMs with tools for complicated reasoning. While Logic-LM performs better, it still struggles indicating that combining LLMs with symbolic solvers is not enough for structured reasoning problems.
Our results indicate that few-shot prompting and Logic-LM are unable to leverage the structure of fcore problems. As problem instance size grows, Logic-LM tends to make syntactic mistakes in the generated constraints with smaller LLMs (Table 2). With larger LLMs syntax mistakes in the constraints reduce, but they still remain semantically incorrect and could not be corrected through feedback. In contrast, Puzzle-LM exploits the first order structure of these problems by generating instance-agnostic programs during training that can then be corrected through feedback from solved examples. They can then be used independently at inference to solve any input instance, highlighting the importance of using program interpreters along with symbolic solvers.
@@ 7.1 Puzzle-LM vs Baselines
Correct Output
7.41% 39.91%
Incorrect Output
46.32% 37.90%
Inefficient Solution
2.90% 2.90%
Syntactic Errors
25.62% 6.36%
Errors in Converting Symbolic Output to Text
17.73% 12.93%
Table 2: Logic-LM’s performance on PuzzleBench evaluated with feedback to fix syntactic mistakes.
Table 3 compares the effect of feedback and multiple runs on PAL and Puzzle-LM. Puzzle-LM outperforms PAL by $27.1\%$ on PuzzleBench (with refinement). This highlights the benefits of offloading the reasoning to a symbolic solver. When LLMs are forced to perform complicated reasoning while writing programs, they tend to produce brute-force solutions which are either often incorrect or extremely slow (see Table-6 in the appendix). Puzzle-LM tends to be much more efficient in terms of inference-time given that the solver can exploit the nature of the input instance while performing the reasoning unlike brute-force programs. Interestingly, feedback from solved examples and re-runs are more effective in improving performance for Puzzle-LM as indicated by the larger gains with increasing number of feedback rounds and runs. We attribute this to the fact that the programs generated by Puzzle-LM are declarative in nature, while PAL produces imperative programs which are difficult to correct.
7.2 Effect of Feedback on Solved Examples
Figure 4 describes the effect of multiple rounds of feedback for Puzzle-LM. Feedback helps performance significantly; utilizing 4 rounds of feedback improves reasoning abilities by $25.36\%$. Even the largest LLMs commit errors, making it important to verify and correct their work. But feedback on its own is not enough, a single run might end-up in a wrong reasoning path, which is not corrected by feedback making it important to utilize multiple runs for effective reasoning. Utilizing 5 runs improves the performance by additional $12.32\%$ after which the gains tend to saturate. Performance also increases with an increase in the number of solved examples (Table 4). Each solved example helps in detecting and correcting different errors. However, performance saturates at 7 solved examples and no new errors are discovered, even with additional training data.
@@ 7.2 Effect of Feedback on Solved Examples
PAL 48.51% 54.52% 57.83% 58.98% 59.80%
Puzzle-LM 60.48% 71.69% 77.37% 80.38% 80.98%
$\uparrow$ 11.97 % $\uparrow$ 17.17 % $\uparrow$ 19.99 % $\uparrow$ 21.4 % $\uparrow$ 21.18 %
(a) Effect of number of rounds of feedback for a single run
@@
PAL 59.80% 62.45% 63.93% 65.23% 66.45&
Puzzle-LM 80.98% 86.75% 89.88% 92.08% 93.55%
$\uparrow$ 21.18 % $\uparrow$ 24.3 % $\uparrow$ 25.95 % $\uparrow$ 26.85 % $\uparrow$ 27.1 %
(b) Effect of number of runs each having 4 rounds of feedback
Table 3: Comparative analysis between PAL and Puzzle-LM on PuzzleBench for GPT-4-Turbo
Number of Feedback Rounds
0 1 2 3 4
N = 0
60.48% 60.48% 60.48% 60.48% 60.48%
N = 1
60.48% 69.47% 73.99% 75.73% 75.75%
N = 4
60.48% 71.43% 76.87% 79.52% 79.54%
N = 7
60.48% 71.69% 77.37% 80.38% 80.98%
N = 10
60.48% 71.69% 77.37% 80.38% 80.98%
Table 4: Effect of number of solved examples (N) for feedback and number of feedback rounds for Puzzle-LM with GPT-4-Turbo on PuzzleBench for a single run.
@@
Figure 5: Effect of increasing problem instance size on baselines and Puzzle-LM for GPT-4-Turbo.
7.3 Effect of Problem Instance Size
@@
We compare performance of Puzzle-LM and other baselines against varying problem instance sizes (see Figure 5) for 3 problems in PuzzleBench (sudoku, sujiko and magic-square). Increasing the problem instance size increases the number of variables, accompanying constraints and reasoning steps required to reach the solution. We observe that the performance of Puzzle-LM and PAL, which are programmatic approaches, is relatively robust against increase in size of input instances. In comparison, performance of Logic-LM and few-shot prompting declines sharply. Performance of PAL declines in some cases because of inefficient programs generated, which are unable to find a solution in the time limit.
8 Qualitative Analysis
We now discuss our insights on where Puzzle-LM excels and where the largest LLMs still struggle. Qualitatively, the problems in PuzzleBench can be broadly categorized into three categories:

### v3

@@ 5.1 Our Approach
Figure 3: SymPro-LM: Solid lines indicate the main flow and dotted lines indicate feedback pathways.
@@
  • Prompting LLMs: The LLM is prompted with $NL(\mathcal{C})$, $NL(\mathcal{X})$, $NL(\mathcal{Y})$ (see Figure 2) to generate an input-agnostic program $\psi$. The LLM is instructed to write $\psi$ to read an input from a file, convert it to a symbolic representation according to the inferred specification of the problem, pass the symbolic representation to the solver and then use the solution from the solver to generate the output in the desired format. The LLM is also prompted with information about the solver and its underlying language. Optionally we can also provide the LLM with a subset of $\mathcal{D}_{\mathcal{P}}$ (see Appendix B.3 for exact prompts).
@@
  • Symbolic Solver: $\psi$ can convert any input instance $x$ to $(E_{x},V_{x})$ which it passes to the symbolic solver. The solver is agnostic to how the representation $(E_{x},V_{x})$ was created and tries to find an assignment $A_{x}$ to $V_{x}$ which satisfies $E_{x}$ which is passed back to $\psi$ (see Appendix E.1 for sample programs generated).
@@
  • Generating the Final Output: $\psi$ then uses $\mathcal{A}_{x}$ to generate the predicted output $\hat{y}$. This step is need because the symbolic representation was created by $\psi$ and it must recover the desired output representation from $\mathcal{A}_{x}$, which might not be straightforward for all problem representations.
@@
Refinement via Solved Examples: We make use of $\mathcal{D}_{\mathcal{P}}$ to verify and (if needed) make corrections to $\psi$. For each $(x,y)\in\mathcal{D}_{\mathcal{P}}$ (solved input-output pair), we run $\psi$ on $x$ to generate the prediction $\hat{y}$, during which the following can happen: 1) Errors during execution of $\psi$; 2) The solver is unable to find $\mathcal{A}_{x}$ under a certain time limit; 3) $\hat{y}\neq y$, i.e. the predicted output is incorrect; 4) $\hat{y}=y$, i.e. the predicted output is correct. If for any training input one of the first three cases occur we provide automated feedback to the LLM through prompts to improve and generate a new program. This process is repeated till all training examples are solved correctly or till a maximum number of feedback rounds is reached. The feedback is simple in nature and includes the nature of the error, the actual error from the interpreter/symbolic solver and the input instance on which the error was generated. For example, in the case where the output doesn’t match the gold output we prompt the LLM with the solved example it got wrong and the expected solution. Appendix B contains details of feedback prompts.
@@
It is possible that a single run of SymPro-LM (along with feedback) is unable to generate the correct solution for all training examples – so, we restart SymPro-LM multiple times for a given problem. Given the probabilistic nature of LLMs a new program is generated at each restart and a new feedback process continues. For the final program, we pick the best program generated during these runs, as judged by the accuracy on the training set. Figure  3 describes our entire approach diagrammatically.
SymPro-LM for Non-First Order Reasoning Datasets: For datasets that are not first-order in nature, a single program does not exist which can solve all problems, hence we prompt the LLM to generate a new program for each test set instance. Thus we cannot use feedback from solved examples and we only use feedback to correct syntactic mistakes (if any). The prompt contains an instruction to write a program which will use a symbolic solver to solve the problem. Additionally, we provide details about the solver to be used. The prompt also contains in-context examples demonstrating sample programs for other logical reasoning questions. The LLM should parse the logical reasoning question and extract the corresponding facts/rules which it needs to pass to the solver (via the program). Once the solver returns with an answer, it is passed back to the program to generate the final output.
@@ 6 Experimental Setup
Our experiments answer these research questions. (1) How does SymPro-LM compare with other LLM-based reasoning approaches on fcore problems? (2) How useful is using feedback from solved examples and multiple runs for fcore problems? (3) How does SymPro-LM compare with other methods on other existing (non-first order) logical reasoning benchmarks? (4) What is the nature of errors made by SymPro-LM and other baselines?
@@
Baselines: On FCoReBench, we compare our method with 4 baselines: 1) Standard LLM prompting, which leverages in-context learning to directly answer the questions; 2) Program-aided Language Models, which use imperative programs for reasoning and offload the solution step to a program interpreter; 3) Logic-LM, which offloads the reasoning to a symbolic solver. 4) Tree-of-Thoughts (ToT) Yao et al. (2023), which is a search based prompting technique. These techniques (Yao et al., 2023; Hao et al., 2023) involve considerable manual effort for writing specialized prompts for each problem and are estimated to be 2-3 orders of magnitude more expensive than other baselines. We thus decide to present a separate comparison with ToT on a subset of FCoReBench (see Appendix C.1.1 for more details regarding ToT experiments). We use Z3 (De Moura and Bjørner, 2008) an efficient SMT solver for experiments with Logic-LM and SymPro-LM. We use the Python interpreter for experiments with PAL and SymPro-LM. We also evaluate refinement for PAL and SymPro-LM by using 5 runs each with 4 rounds of feedback on solved examples for each problem. We evaluate refinement for Logic-LM by providing 4 rounds of feedback to correct syntactic errors in constraints (if any) for each problem instance. We decide not to evaluate SAT-LM given its conceptual similarity to Logic-LM having being proposed concurrently.
@@
Models: We experiment with 3 LLMs: GPT-4-Turbo (gpt-4-0125-preview) (OpenAI, 2023) which is a SOTA LLM by OpenAI, GPT-3.5-Turbo (gpt-3.5-turbo-0125), a relatively smaller LLM by OpenAI and Mixtral 8x7B (open-mixtral-8x7b) (Jiang et al., 2024), an open-source mixture-of-experts model developed by Mistral AI. We set the temperature to 00 for few-shot prompting and Logic-LM for reproducibility and to $0.7$ to sample several runs for PAL and SymPro-LM.
@@
Prompting LLMs: Each method’s prompt includes the natural language description of the problem’s rules and the input-output format, along with two solved examples. No additional intermediate supervision (e.g., SMT or Python program) is given in the prompt. For few-shot prompting we directly prompt the LLM to solve each test set instance separately. For PAL we prompt the LLM to write an input-agnostic Python program which reads the input from a file, reasons to solve the input and then writes the solution to another file, the program generated is run on each testing set instance. For Logic-LM for each test set instance we prompt the LLM to convert it into its symbolic representation which is then fed to a symbolic solver, the prompt additionally contains the description of the language of the solver. We then prompt the LLM with the solution from the solver and ask it to generate the output in the desired format (see Section 5). Prompt templates are detailed in Appendix B and other experimental details can be found in Appendix C.
@@
Metrics: For each problem, we use the associated verification script to check the correctness of the candidate solution for each test instance. This script computes the accuracy as the fraction of test instances solved correctly, using binary marking assigning 1 to correct solutions and 0 for incorrect ones. We report the macro-average of test set accuracies across all problems in FCoReBench.
Additional Datasets: Apart from FCoReBench, we also evaluate SymPro-LM on 3 additional logical reasoning datasets: (1) LogicalDeduction from the BigBench (bench authors, 2023) benchmark, (2) ProofWriter (Tafjord et al., 2021) and (3) PrOntoQA (Saparov and He, 2023a). In addition to other baselines, we also compare with Chain-of-Thought (CoT) prompting (Wei et al., 2022c), as it performs significantly better than standard prompting for such datasets. Recall that these benchmarks are not first-order in nature i.e. each problem is accompanied with a single instance (despite the rules potentially being first-order) and hence we have to run SymPro-LM (and other methods) separately for each test instance (see Appendix C.2 for more details).
@@ 7 Results
Table 1 describes the main results for FCoReBench. Unsurprisingly, GPT-4-Turbo is hugely better than other LLMs. Mixtral 8x7B struggles on our benchmark indicating that smaller LLMs (even with mixture of experts) are not as effective at complex reasoning. Mixtral in general does badly, often doing worse than random (especially when used without refinement). PAL and SymPro-LM tend to perform better than other baselines benefiting from the vast pre-training of LLMs on code (Chen et al., 2021). Logic-LM performs rather poorly with smaller LLMs indicating that they struggle to invoke symbolic solvers directly.
Table 1: Results for FCoReBench. - / + indicate before / after refinement. Performance for random guessing is 20.13%.
Few-Shot PAL Logic-LM SymPro-LM
Model Prompting - + - + - +
Mixtral 8x7B 25.06% 14.98% 36.09% 0.21% 2.04% 8.08% 30.09%
GPT-3.5-Turbo 27.02% 32.66% 49.19% 6.04% 6.58% 17.08% 50.35%
GPT-4-Turbo 29.33% 47.42% 66.40% 34.11% 38.51% 50.94% 83.37%
Hereafter, we focus primarily on GPT-4-Turbo’s performance, since it is far superior to other models. SymPro-LM outperforms few-shot prompting and Logic-LM across all problems in FCoReBench. On average the improvements are by an impressive $54.04\%$ against few-shot prompting and by $44.86\%$ against Logic-LM (with refinement). Few-shot prompting solve less than a third of the problems with GPT-4-Turbo, suggesting that even the largest LLMs cannot directly perform complex reasoning. While Logic-LM performs better, it still isn’t that good either, indicating that combining LLMs with symbolic solvers is not enough for such reasoning problems.
Table 2: Logic-LM’s performance on FCoReBench evaluated with refinement.
@@
Correct Output 6.58% 38.51%
Incorrect Output 62.11% 52.06%
Timeout Error 2.375% 2.49%
Syntactic Error 29.04% 6.91%
Table 3: Error analysis at a program level for GPT-4-Turbo before and after refinement for PAL and SymPro-LM. Results are averaged over all runs for a problem and further over all problems in FCoReBench.
Outcome PAL SymPro-LM
(Before / After) (Before / After)
Incorrect Program 70% / 57% 58% / 38%
Semantically Incorrect Program 62% / 49.5% 29% / 20.5%
Python Runtime Error 7% / 4.5% 13.5% / 5.5%
Timeout 1% / 3% 15.5% / 12%
Further qualitative analysis suggests that Logic-LM gets confused in handling the structure of fcore problems. As problem instance size grows, it tends to make syntactic mistakes with smaller LLMs (Table 3). With larger LLMs, syntactic mistakes reduce, but constraints still remain semantically incorrect and do not get corrected through feedback.
Often this is because LLMs are error-prone when enumerating combinatorial constraints, i.e., they struggle with executing implicit for-loops and conditionals (see Appendix F). In contrast, SymPro-LM and PAL manage first order structures well, since writing code for a loop/conditional is not that hard, and the correct loop-execution is done by a program interpreter. These (size-invariant) programs then get used independently without any LLM call at inference time to solve any input instance – easily generalizing to larger instances – highlighting the benefit of using a program interpreter for such combinatorial problems.
At the same time, PAL is also not as effective on FCoReBench. Table 4 compares the effect of feedback and multiple runs on PAL and SymPro-LM. SymPro-LM outperforms PAL by $16.97\%$ on FCoReBench (with refinement). When LLMs are forced to write programs for performing complicated reasoning, they tend to produce brute-force solutions that often are either incorrect or slow (see Table-8 in the appendix). This highlights the value of offloading reasoning to a symbolic solver. Interestingly, feedback from solved examples and re-runs is more effective (Table  3) for SymPro-LM, as also shown by larger gains with increasing number of feedback rounds and runs (Table  4). We hypothesize that this is because declarative programs (generated by SymPro-LM) are easier to correct, than imperative programs (produced by PAL).
Table 4: Comparative analysis between PAL and SymPro-LM on FCoReBench for GPT-4-Turbo.
@@
PAL 47.42% 54.00% 57.09% 58.82% 59.92%
SymPro-LM 50.94% 62.54% 68.52% 71.12% 71.96%
$\uparrow$3.52$\uparrow$8.54$\uparrow$11.43$\uparrow$12.3$\uparrow$12.04%
(a) Effect of feedback rounds for a single run
@@
PAL 59.92% 62.54% 63.95% 65.19% 66.40%
SymPro-LM 71.96% 77.21% 80.06% 82.06% 83.37%
$\uparrow$12.04$\uparrow$14.67$\uparrow$16.11$\uparrow$16.87$\uparrow$16.97%
(b) Effect of multiple runs each with 4 feedback rounds
Table 5: Accuracy and cost comparison between ToT prompting and SymPro-LM with GPT-4-Turbo for 3 problems in FCoReBench. Costs are per test instance for ToT and one time costs per problem for SymPro-LM.
Problem Instance size ToT prompting SymPro-LM
Accuracy Cost Accuracy Cost
Latin Squares 3x3 46.33% \$0.1235 100% \$0.02
4x4 32.5% \$0.5135 100% \$0.02
Magic Square 3x3 26.25% \$0.4325 100% \$0.02
4x4 8% \$0.881 100% \$0.02
Sujiko 3x3 7.5% \$0.572 100% \$0.02
4x4 0% \$1.676 100% \$0.02
Comparison with ToT Prompting: Table 5 compares SymPro-LM with ToT prompting on 3 problems. SymPro-LM is far superior in terms of cost and accuracy, indicating that even the largest LLMs cannot do complex reasoning on problems with large search depths and branching factors, despite being called multiple times with search-based prompting. Due to its programmatic nature, SymPro-LM generalizes even better to larger instances and is also hugely cost effective, as there is no need to call an LLM for each instance separately. We do not perform further experiments with ToT prompting, due to cost considerations.
@@
Figure 4: Effect of increasing problem instance size on baselines and SymPro-LM for GPT-4-Turbo.
@@
Effect of Problem Instance Size: We now report performance of SymPro-LM and other baselines against varying problem instance sizes (see Figure 4) for 3 problems in FCoReBench (sudoku, sujiko and magic-square). Increasing the problem instance size increases the number of variables, accompanying constraints and reasoning steps required to reach the solution. We observe that being programmatic SymPro-LM and PAL, are relatively robust against increase in size of input instances. In comparison, performance of Logic-LM and few-shot prompting declines sharply. PAL programs are often inefficient and may see performance drop when they fail to find a solution within the time limit.
(a) Effect of feedback
(b) Effect of multiple runs
(c) Effect of # of solved examples
Figure 5: Effect of feedback and multiple runs with GPT-4-Turbo. (a) and (b) show results with 10 solved examples for feedback where dashed lines show results for individual problems in FCoReBench, with coloured lines highlighting specific problems and the red bold line represents the average effect across all problems. (c) shows the effect of number of solved examples used for feedback in a single run.
Effect of Feedback on Solved Examples: Figure 5(a) describes the effect of multiple rounds of feedback for SymPro-LM. Feedback helps performance significantly; utilizing 4 feedback rounds improves performance by $21.02\%$. Even the largest LLMs commit errors, making it important to verify and correct their work. But feedback on its own is not enough, a single run might end-up in a wrong reasoning path, which is not corrected by feedback making it important to utilize multiple runs for effective reasoning. Utilizing 5 runs improves the performance by additional $11.41\%$ (Figure 5(b)) after which the gains tend to saturate. Performance also increases with an increase in the number of solved examples (Figure 5(c)). Each solved example helps in detecting and correcting different errors. However, performance tends to saturate at 7 solved examples and no new errors are discovered/corrected, even with additional training data.
7.1 Results on Other Datasets
Table 6: Results for baselines & SymPro-LM on other benchmarks. Best results with each LLM are highlighted.
GPT-3.5-Turbo-0125 GPT-4-Turbo-0125
Dataset Direct CoT PAL Logic-LM SymPro-LM Direct CoT PAL Logic-LM SymPro-LM
Logical Deduction 39.66 % 50.66 % 66.33 % 71.00 % 78.00 % 65.33 % 76.00 % 81.66 % 82.67 % 94.00 %
ProofWriter 40.50 % 57.16 % 50.5 % 70.16 % 74.167 % 46.5 % 61.66 % 76.29 % 74.83 % 89.83 %
PrOntoQA 49.60 % 83.20 % 98.40 % 72.20 % 97.40 % 83.00 % 98.80 % 99.80 % 91.20 % 97.80 %
Table 6 reports the performance on non-first order datasets. SymPro-LM outperforms all other baselines on ProofWriter and LogicalDeduction, particularly Logic-LM. This showcases the value of integrating LLMs with symbolic solvers through programs, even for standard reasoning tasks. These experiments suggest that LLMs translate natural language questions into programs using solvers much more effectively than into symbolic formulations directly. We attribute this to the vast pre-training of LLMs on code (Brown et al., 2020; Chen et al., 2021). For instance, on the LogicalDeduction benchmark, while Logic-LM does not make syntactic errors during translation it often makes logical errors. These errors significantly decrease when LLMs are prompted to produce programs instead (Figure 6(b)). Error analysis on ProofWriter and PrOntoQA reveals that for more complex natural language questions, LLMs also start making syntactic errors during translation as the number of rules/facts start increasing. With SymPro-LM these errors are vastly reduced because, apart from the benefit from pre-training, LLMs also start utilizing programming constructs like dictionaries and loops to make most out of the structure in these problems (Figure 6(a)). PAL and CoT perform marginally better on PrOntoQA because the reasoning style for problems in this dataset involves forward-chain reasoning which aligns with PAL’s and CoT’s style of reasoning. Integrating symbolic solvers is not as useful for this dataset, but still achieves competitive performance.
(a) PrOntoQA
(b) LogicalDeduction
Figure 6: Examples highlighting benefits of integrating LLMs with symbolic solver through programs.
8 Discussion
We analyze FCoReBench to identify where LLMs excel and where the largest models still struggle. Based on SymPro-LM’s performance, we categorize FCoReBench problems into three broad groups.